A rectangle is 13 inches longer than it is wide and the perimeter is 226 inches. What is the length and width?
Answer by rfer(16322)

width=x=50
length=x+13=63
-----------------
2(x+13)+2x=226
2x+26+2x=226
4x=200
x=50
---------------
2*50+2*63=226